# Nord-Parl-TTS: Finnish and Swedish TTS Dataset from Parliament Speech

### v1

@@ 1 Introduction
Recent work on building extensive and diverse datasets from speech found in the wild [6, 7, 8, 9] directly targets this challenge. In particular, Emilia [6] proposed a unified data processing pipeline and released 101k hours of multilingual speech data. This was followed by several works [7, 8, 9] that further extended the data quality and quantity. While these efforts mainly target high-resource languages like English and Mandarin Chinese, languages with lower resources are often overlooked. As a step toward addressing this imbalance in the Nordic region, we present Nord-Parl-TTS, an open in-the-wild TTS dataset for Finnish and Swedish. We adapted the Emilia data processing pipeline and extracted ready-to-use TTS data from parliament speech recordings, resulting in 900 hours of Finnish and 5090 hours of Swedish TTS data. To further support the development of TTS for these languages, we propose unified evaluation sets following [1, 10]. Finally, we conduct benchmark experiments on two open-source TTS systems to demonstrate their applicability for model development and evaluation.
@@ 2 Related Work
Finnish parliament recordings have been transformed into the Finnish Parliament ASR corpus [14]. However, further transforming the ASR corpus into a TTS dataset is not an optimal solution for several reasons. First, the transcripts are corrected for hesitations, repetitions, and slips of the tongue. They are also edited to replace spoken and spontaneous language with equivalent written language for clarity [14]. Second, the ASR corpus is segmented based on the word boundary, meaning that an utterance can start and end in the middle of a sentence. The corrected transcripts make it difficult for TTS models to learn the right pronunciations, and the word-level segmentation negatively affects the sentence-level prosody modeling.
@@ 3.2 Swedish
We process the RixVox-v2 for the Swedish dataset. The pipeline is shown in the yellow arrows in Figure 1. We first resample the audio to 24 kHz and use the same speech separation model in Section 3.1 to clean out non-speech tracks. Then we cut the utterance using the existing sentence-level timestamp from the dataset if the gap between two sentences is longer than 2 seconds. The same speaker diarization model in Section 3.1 is used to exclude segmented utterances with multiple speakers. To evaluate the intelligibility of the utterance, we transcribe the utterance using the Swedish Whisper-large [19] and filter out those utterances that have a Word Error Rate (WER) larger than 10%. We allow a WER of 10% because the sentence-level transcripts and timestamps of RixVox-v2 are also derived from ASR models, but they are not perfectly accurate. On the other hand, Swedish Whisper-large is trained on many large-scale Swedish datasets including RixVox-v2, making it more accurate than the original transcripts [19]. Moreover, we wish to filter out unintelligible speech while keeping the scale of the dataset as large as possible. Finally, we filter out utterances with a DNSMOS P.835 OVRL score less than 3.0. This yields a 5090-hour Swedish dataset. The DNSMOS score distribution and the duration distribution of this dataset are shown in the yellow pillars in Figure 2(a) and Figure 2(b), respectively. We notice that our pipeline for Swedish results in a monotonically increasing duration distribution in the dataset, which is different from existing datasets that are dominated by utterances in 5 to 15 seconds in duration.
@@ 4 Experiments
For subjective evaluations, we evaluate the human-likeness of the synthetic speech using Comparative Mean Opinion Score (CMOS) and speaker similarity using Speaker Mean Opinion Score (SMOS). For CMOS, we ask the participants whether sample A or sample B is more human-like, with a score between -3 to 3 and an increment of 1, where -3 means sample A is definitely more human-like, 0 is unsure, and 3 means sample B is definitely more human-like. One of the two samples is the ground truth speech, and the other is the synthetic speech. For SMOS, we ask how sample B (synthetic speech) sounds like the same speaker as sample A (target speech), and the participants should rate between -2 and 2 with an increment of 1, where -2 represents definitely not the same speaker and 2 represents definitely the same speaker. We map SMOS back to a 1 to 5 scale when calculating the statistics. For both languages, we insert several language skill checks into each questionnaire, which instruct the participants to rate certain scores on certain questions in the testing language. Our participants are recruited via Prolific [^7] https://www.prolific.com/, a crowd-sourcing platform. For both languages, we screen the participants by their primary language, and with a history of at least 98% acceptance rate. For Finnish, we also recruit students and employees from the university to ensure the desired number of participants. We exclude the submissions that fail the language skill checks from the final results.

### v2

@@ 1 Introduction
Recent work on building extensive and diverse datasets from speech found in the wild [8, 17, 16, 11] directly targets this challenge. In particular, Emilia [8] proposed a unified data processing pipeline and released 101k hours of multilingual speech data. This was followed by several works [17, 16, 11] that further extended the data quality and quantity. While these efforts mainly target high-resource languages like English and Mandarin Chinese, languages with lower resources are often overlooked. As a step toward addressing this imbalance in the Nordic region, we present Nord-Parl-TTS, an open in-the-wild TTS dataset for Finnish and Swedish. We adapted the Emilia data processing pipeline and extracted ready-to-use TTS data from parliament speech recordings, resulting in 900 hours of Finnish and 5090 hours of Swedish TTS data. We choose parliament speech as the source because it is open-access and free to process and distribute as long as the original content is not modified. To further support the development of TTS for these languages, we propose unified evaluation sets following [5, 1]. Finally, we conduct benchmark experiments on two open-source TTS systems to demonstrate their applicability for model development and evaluation.
@@ 2 Related Work
Finnish parliament recordings, which include nearly 3000 hours of parliament sessions between 2015 and 2020, have been transformed into the Finnish Parliament ASR corpus [25]. However, further transforming the ASR corpus into a TTS dataset is not an optimal solution for several reasons. First, the transcripts are corrected for hesitations, repetitions, and slips of the tongue. They are also edited to replace spoken and spontaneous language with equivalent written language for clarity [25]. Second, the ASR corpus is segmented based on the word boundary, meaning that an utterance can start and end in the middle of a sentence. The corrected transcripts make it difficult for TTS models to learn the right pronunciations, and the word-level segmentation negatively affects the sentence-level prosody modeling.
@@ 3.2 Swedish
We process the RixVox-v2 for the Swedish dataset. The pipeline is shown in the yellow arrows in Figure 1. We first resample the audio to 24 kHz and use the same speech separation model in Section 3.1 to clean out non-speech tracks. Then we cut the utterance using the existing sentence-level timestamp from the dataset if the gap between two sentences is longer than 2 seconds. The same speaker diarization model in Section 3.1 is used to exclude segmented utterances with multiple speakers. To evaluate the intelligibility of the utterance, we transcribe the utterance using the Swedish Whisper-large [24] and filter out those utterances that have a Word Error Rate (WER) larger than 10%. We allow a WER of 10% because the sentence-level transcripts and timestamps in RixVox-v2 contain similar corrections to those in the Finnish Parliament ASR corpus [25]. The Swedish Whisper, trained on multiple large-scale Swedish datasets including RixVox-v2 [24], represents the content more faithfully. Moreover, we wish to filter out unintelligible speech while keeping the scale of the dataset as large as possible. Finally, we filter out utterances with a DNSMOS P.835 OVRL score less than 3.0. This yields a 5090-hour Swedish dataset. The DNSMOS score distribution and the duration distribution of this dataset are shown in the yellow pillars in Figure 2(a) and Figure 2(b), respectively. We notice that our pipeline for Swedish results in a monotonically increasing duration distribution in the dataset, which is different from existing datasets that are dominated by utterances in 5 to 15 seconds in duration.
@@ 4 Experiments
For subjective evaluations, we evaluate the human-likeness of the synthetic speech using Comparative Mean Opinion Score (CMOS) and speaker similarity using Speaker Mean Opinion Score (SMOS). We use CMOS instead of “standard” Mean Opinion Score (MOS) to avoid perpetuating problematic practices [14]. For CMOS, we ask the participants whether sample A or sample B is more human-like, with a score between -3 to 3 and an increment of 1, where -3 means sample A is definitely more human-like, 0 is unsure, and 3 means sample B is definitely more human-like. One of the two samples is the ground truth speech, and the other is the synthetic speech. For SMOS, we ask how sample B (synthetic speech) sounds like the same speaker as sample A (target speech), and the participants should rate between -2 and 2 with an increment of 1, where -2 represents definitely not the same speaker and 2 represents definitely the same speaker. We map SMOS back to a 1 to 5 scale when calculating the statistics. For both languages, we insert several language skill checks into each questionnaire, which instruct the participants to rate certain scores on certain questions in the testing language. Our participants are recruited via Prolific [^7] https://www.prolific.com/, a crowd-sourcing platform. For both languages, we screen the participants by their primary language, and with a history of at least 98% acceptance rate. For Finnish, we also recruit students and employees from the university to ensure the desired number of participants. We exclude the submissions that fail the language skill checks from the final results.
5 Results